Suppose the prices of the commodities A and B are p1 and p2 respectively. The consumer purchases X1 unit of A and X2 units of B. Suppose, the consumer has an income denoted by M and the consumer would spend the entire amount of M on these two commodities. How are p1, p2, X1, X2 and M related?


1
Expert's answer
2021-08-03T11:51:08-0400

We are given the price of commodity A is p1 and the price of commodity B is p2.

Now, X1 units of commodity A is purchased and X2 units of the commodity B is purchased 

1 unit of commodity A costs p1

This means X1 units of commodity A costs X1p1

Similarly, 

1 unit of commodity B costs p2

This means X2 units of commodity B costs X2p2


Now, the total income is M and the entire income is spent on purchasing the two commodities.

Therefore, the sum of the price of the two commodities is M.

This means:

"X_1p_1+X_2p_2=M"

Therefore, "p_1, p_2, X_1, X_2, M" are related by the equation "X_1p_1+X_2p_2=M"

Answer: "p_1, p_2, X_1, X_2, M" are related by the equation "X_1p_1+X_2p_2=M"